OpenStudy (anonymous):

A rocket is launched from ground level with an initial vertical velocity (v0) of 176 ft/s. After how many seconds will the rocket hit the ground? (Hint: h(t) = -16t2 + v0t + h0). __________ s. Enter a number only.

sam (.sam.):

v=u+at v=0 at max height, 0=174+(-9.81)t t=17.7s Return trip= 17.7 x 2=35.4sec